Solve the system of equations
x + y = 5
x - y = -9

Answer by ptth176(50)

x + y = 5(1)
x - y = -9(2)
(1)+(2)=>2x=-4>x=-2
(1)=>-2+y=5=>y=7
Finally:x=-2;y=7